OpenStudy (anonymous):

The function f(x) = 2x^2 + 3x + 5, when evaluated, gives a value of 19. What is the function’s input value? A. 1 b. -1 C. 2 D. -2 E. -3

OpenStudy (campbell_st):

have you substituted any values to check them...?

OpenStudy (campbell_st):

as an example x= 1 y = 2(1)^2 + 3(1) + 5 = 10 so its not x = 1

OpenStudy (anonymous):

o that's all i have to do?

OpenStudy (saiyan):

yep

OpenStudy (anonymous):

@campbell_st im lost i thought i had it but i dont

OpenStudy (campbell_st):

ok so the next value is x = -1 substitute that y = 2(-1)^2 + 3(-1) + 5 you need a little care here as squaring a negative gives a positive so y = 2 - 3 + 5 y = 4 so it't not x = -1 just repeat the process

OpenStudy (anonymous):

ok thanks